OpenStudy (anonymous):

Find functions f and g so that h(x)=(f*g)(x) -h(x)=I 6x+8 I & -h(x)=9/square root 4x+3

myininaya (myininaya):

so you are looking for two factors such that their product is |6x+8|? Or is that * suppose to be meant as a composition? in which case you are looking for a function where you can compose another function and get |6x+8|

myininaya (myininaya):

Example say you have \[h(x)=f(g(x)) \\ \text{ and you are given } h(x)=(9x)^2 \\ \text{ then the outside function is } x^2 \\ \text{ where the inside function is } 9x \\ \\ \text{ So } f(x)=x^2\ \ \text{ and } g(x)=9x \\ \text{ also people normally use a small little circle to } \\ \text{ mean composition } \\ \text{ like } (f \circ g)(x)=f(g(x))\]